# SCORE: Story Coherence and Retrieval Enhancement for AI Narratives

### v1

@@ Abstract
Large Language Models (LLMs) excel at generating creative narratives but struggle with long-term coherence and emotional consistency in complex stories. To address this, we propose SCORE (Story Coherence and Retrieval Enhancement), a framework integrating three components: 1) Dynamic State Tracking (monitoring objects/characters via symbolic logic), 2) Context-Aware Summarization (hierarchical episode summaries for temporal progression), and 3) Hybrid Retrieval (combining TF-IDF keyword relevance with cosine similarity-based semantic embeddings). The system employs a temporally-aligned Retrieval-Augmented Generation (RAG) pipeline to validate contextual consistency. Evaluations show SCORE achieves 23.6% higher coherence (NCI-2.0 benchmark), 89.7% emotional consistency (EASM metric), and 41.8% fewer hallucinations versus baseline GPT models. Its modular design supports incremental knowledge graph construction for persistent story memory and multi-LLM backend compatibility, offering an explainable solution for industrial-scale narrative systems requiring long-term consistency.
@@ 1 Introduction
Figure 1: A conceptual overview of common issues in LLM-generated stories (left), why existing methods struggle (middle), and how our proposed SCORE framework addresses these problems through dynamic state tracking, context-aware summarization, and hybrid retrieval (right).
Large Language Models (LLMs) have demonstrated significant capabilities in generating long-form narratives, such as serialized stories or novels, by leveraging large-scale architectures and vast amounts of training data Tao et al. (2024). However, maintaining narrative consistency over extended texts, especially in terms of character development and emotional coherence, remains a major challenge McAdams (2006). For instance, Khatun & Brown (2024) pointed out that achieving thematic consistency and managing dynamic plot states is crucial for maintaining the logical flow of a story. In practice, LLMs often show inconsistencies when characters or key plot items reappear without proper justification, disrupting the overall narrative structure.
Likewise, Liu et al. (2024) highlight the difficulties in managing multimodal elements within long-form narratives, noting that inconsistencies in character behavior or emotional tone can negatively impact reader engagement. These observations indicate a need for more structured methods of narrative generation that can better handle character arcs, plot events, and emotional progression over the course of a story.
In addition, recent works have highlighted the importance of memory mechanisms in LLM-based agentsLi et al. (2024); Liang et al. (2024). Zhang et al. (2024) present a comprehensive overview of these mechanisms, illustrating their role in reducing narrative inconsistencies that occur in interactive tasks and story generation. Additionally, Park et al. (2023) introduced generative agents that simulate human-like behavior using memory modules. These agents track the state of a wide array of interactable objects to ensure coherent reasoning and stable interactions within a simulated society. Their findings informed our own framework design.
In this work, we build upon recent advancements in Retrieval-Augmented Generation (RAG) Lewis et al. (2020), which dynamically incorporates relevant external context to enhance narrative coherence. Building on these developments, we propose SCORE, a framework designed to evaluate three critical aspects of long-form narrative generation: character consistency, emotional coherence, and logical tracking of key plot elements. Our key contributions are:
@@
  • We surpass the baseline GPT model Radford et al. (2019) in detecting continuity errors and preserving overall narrative coherence.
2 Related Work
Foundations of Storytelling and Interactive Narratives
Storytelling has long been recognized as a fundamental means of human communication, capable of conveying complex ideas and deep emotional experiences (Rowcliffe, 2004; Miller & Pennycuff, 2008). Early work in interactive narrative planning demonstrated that extended stories could be generated from simple narrative seeds (Porteous et al., 2021). Multi-agent storytelling systems, such as SABRE, further explored the coordination of multiple characters within a shared narrative space to enhance story complexity and realism (Ware & Siler, 2021). These foundational studies have paved the way for more advanced narrative generation methods.
Advancements in AI-Driven Story Generation
Recent advances in natural language processing (NLP) have led to the emergence of large language models (LLMs) that can generate fluent and creative narratives. Models such as GPT-3 (Brown et al., 2020), LaMDA (Thoppilan et al., 2022), PaLM (Chowdhery et al., 2023), and LLaMA (Touvron et al., 2023) have demonstrated impressive generative capabilities (Clark et al., 2021). However, these models often struggle with maintaining long-term coherence and consistent character behavior over extended texts. To address these challenges, recent research has integrated planning and memory components with LLMs. For example, Yang et al. (2024) introduced SEED-Story, a method leveraging multimodal LLMs to generate extended narratives that combine text and visuals. Similarly, Wang & Kreminski (2024) proposed a framework that employs Answer Set Programming to guide and diversify story generation.
Retrieval-Augmented Generation and Dynamic State Tracking
Ensuring narrative continuity over long texts requires effective retrieval and state tracking mechanisms. Retrieval-augmented generation (RAG) techniques have been employed to incorporate relevant historical context into the generation process (Lewis et al., 2020). Methods based on TF-IDF (Ramos et al., 2003) and semantic embeddings via FAISS (Douze et al., 2024) allow models to retrieve and utilize pertinent past content, thereby reinforcing narrative consistency. Moreover, dynamic state tracking approaches—which often draw inspiration from psychological models of narrative coherence (Mcadams, 2006)—have been developed to monitor and correct inconsistencies in character actions and object statuses.
Evaluation for AI-Generated Narratives
As the accessibility of LLMs continues to increase, there is a growing need for accurate and scalable methods to evaluate generated stories. Manual evaluation can require considerable human effort (Celikyilmaz et al., 2020), and many automatic metrics do not align well with human preferences (Chhun et al., 2022). Recent work has begun investigating whether LLMs themselves can serve as evaluators by rating and explaining their assessments of generated narratives (Chhun et al., 2024), suggesting a possible direction for more robust automated evaluation in storytelling research.
Figure 2: The SCORE framework for improving AI-generated story coherence. (a) Extracts key item statuses per episode. (b) Produces detailed analysis and summaries of each episode. (c) Uses RAG to answer user queries and resolve narrative inconsistencies.
@@ 3 Method
Our proposed method, SCORE(illustrated in Figure 2), consists of three main components: (1) Dynamic State Tracking, which monitors and corrects inconsistencies in the states of key narrative elements (e.g., characters and objects) using symbolic logic; (2) Context-Aware Summarization, which automatically generates hierarchical episode summaries to capture plot development and emotional progression; and (3) Hybrid Retrieval, which integrates the outputs of the first two components via a Retrieval-Augmented Generation (RAG) pipeline to enforce contextual consistency across episodes.
@@
3.1 Dynamic State Tracking
@@
By extracting key parts of GPT-4’s analysis, we identify instances where an item reappears in later episodes after being marked as lost or destroyed, without any narrative explanation. Let $S_{i}(t)$ represent the state of item $i$ at time $t$, where $S_{i}(t)\in\{\text{active},\text{lost},\text{destroyed}\}$. If item $i$ reappears at time $t_{k}$ with $S_{i}(t_{k})=\text{active}$ after being previously marked as $S_{i}(t_{k-1})\in\{\text{lost},\text{destroyed}\}$, we flag this as a continuity error. To maintain consistency, the state remains $S_{i}(t_{k-1})$, avoiding an incorrect update. This approach systematically corrects discrepancies in item states, ensuring that the story maintains coherent transitions and avoids erroneous state changes. This method aligns with the psychological framework of narrative coherence proposed by  Mcadams (2006), which defines the narrative entropy. By preventing unexplained reactivations of lost or destroyed items, our approach reduces narrative entropy and improves continuity. See section A.1 for details.
3.2 Context-Aware Summarization
@@
For each episode, we produce a detailed evaluation by summarizing the main plot points, character actions, and significant item interactions. Let $A_{c}(t)$ represent the actions of character $c$ at time $t$, and let $I_{i}(t)$ denote the interactions with key item $i$. The model then generates summaries containing $A_{c}(t)$ (character actions), relationships, and changes in emotional states within the episode. It analyzes the specific interactions $I_{i}(t)$ between characters and key items, documenting these for further analysis. This step aggregates relevant narrative information-combining episode summaries, key item interactions $I_{i}(t)$, and character actions $A_{c}(t)$-to facilitate more precise future retrieval. The approach simplifies subsequent analysis of plot and item continuity, reducing redundancy and improving efficiency.
3.3 Hybrid Retrieval
@@
We integrate similarity-based retrieval and sentiment analysis to improve episode evaluation and answer complex queries. It begins by loading summaries, full episode content, and key item states from structured files. The content is segmented into smaller chunks using a text segmenter and embedded into a vector space model using FAISS Douze et al. (2024) and OpenAI embeddings. This vector space enables efficient retrieval of similar episodes for user queries or specific episode analysis.
By integrating Retrieval-Augmented Generation (RAG) with TF-IDF Ramos et al. (2003), the system can more accurately locate and evaluate relevant information within large text collections. RAG introduces an external retrieval module that gathers supporting content before generating final answers. TF-IDF calculates term importance by weighing frequent yet discriminative keywords.
@@
For evaluation, the system retrieves relevant past episodes by calculating cosine similarity scores Rahutomo et al. (2012) between the current episode or query and all other episodes in the vector space. Let $S(e_{c},e_{p})$ represent the similarity score between the current episode $e_{c}$ and a past episode $e_{p}$. The top $N$ episodes with the highest $S(e_{c},e_{p})$ scores are retrieved for further analysis, providing a relevant summary of episodes for evaluation or answering questions.
@@
To further explain the performance gain in complex query answering, especially as shown in Table 2, we provide an information-theoretic justification of how retrieval-augmented generation (RAG) improves narrative coherence following the work of  Tishby et al. (2000). Specifically, we assume that the SCORE mechanism compresses redundant content by retrieving key elements (e.g., characters, objects, emotions), thereby preserving the minimal sufficient statistics of the narrative.
Theorem 3.1 (Optimal Retrieval Distribution via the Information Bottleneck Principle).
and the optimal retrieval probability is
$p(e_{p}|e_{c})=\frac{K(e_{c},e_{p})}{\sum_{e\in\mathcal{E}}K(e_{c},e)},$ (2)
The details and underlying proof are provided in the section A.2.
@@ 4.1 Dataset Preparation
As is shown in Table 5, we constructed a comprehensive story dataset comprising 5,000 episodes from 1,000 GPT-generated stories, categorized into four major genres: science fiction (250 stories), drama (250), fantasy (250), and comedy (250). Each story contains between 10 and 15 episodes, with an average of 12 episodes per story, ensuring a balanced representation of short-, medium-, and long-term narrative arcs. To ensure diversity, we employed model heterogeneity, with outputs from GPT-3.5 (50%) and GPT-4 (50%), utilizing 15 distinct prompt templates per genre, including character-driven, plot-driven, and theme-driven variants. Quality control was implemented through manual filtering by three annotators to exclude stories with less than 80% alignment to the target genres, yielding a Fleiss’ $\kappa$ of 0.78 for inter-annotator agreement. Additionally, complexity stratification was applied, with 30% of stories featuring multi-character interactions (more than four main characters), 20% incorporating non-linear timelines, and 50% containing symbolic object systems (e.g., magical artifacts in fantasy). The stories average 2,000 tokens ($\sigma=450$), with episode lengths ranging from 800 to 5,000 tokens. SCORE was further validated on three standard datasets: NarrativeQA (1,567 human-written stories with annotated plot graphs), BookCorpus (11,038 novels across 16 genres), and WP-STORIES (a human-authored benchmark for coherence evaluation). For cross-dataset evaluation, we aligned genre distributions (SF: 20%, Drama: 30%, Fantasy: 30%, Comedy: 20%) and standardized episode lengths (1,500–2,500 tokens) through truncation and padding. This hybrid approach guarantees the inclusion of both synthetic and human-crafted narratives, ensuring robustness across both industrial-scale and academic-standard benchmarks.
4.2 Metrics
Consistency:Yoo & Cheong (2024) We define this as the fraction of responses that do not conflict with previously stated information. For $N$ responses $r_{1},r_{2},\dots,r_{N}$, let $\mathbf{1}(\mathrm{consistent}(r_{i}))$ be $1$ if $r_{i}$ is consistent and 00 otherwise. Then,
$\mathrm{Cons}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(\mathrm{consistent}(r_{i})).$ (3)
Coherence:Yoo & Cheong (2024) We measure coherence as an average coherence score over $N$ responses. Let $\mathrm{coh}(r_{i})$ be a coherence function returning a numeric value (for instance, based on semantic flow or logical structure). Then,
$\mathrm{Coh}=\frac{1}{N}\sum_{i=1}^{N}\mathrm{coh}(r_{i}).$ (4)
Item Status:Yang & Jin (2024) We define this as the percentage of required items present in the response. For each $r_{i}$, let $\mathrm{req}(r_{i})$ be the set of items that must appear, and $\mathrm{inc}(r_{i})$ be the set of items actually included. Then,
$\mathrm{IS}=\frac{1}{N}\sum_{i=1}^{N}\frac{|\mathrm{req}(r_{i})\cap\mathrm{inc}(r_{i})|}{|\mathrm{req}(r_{i})|}.$ (5)
Complex Question:Yang & Jin (2024) We define this as the fraction of complex questions answered correctly. Let $Q$ be the set of all complex questions posed, and let $\mathbf{1}(\mathrm{correct}(q))$ be $1$ if question $q$ is answered correctly and 00 otherwise. Then,
$\mathrm{CQ}=\frac{1}{|Q|}\sum_{q\in Q}\mathbf{1}(\mathrm{correct}(q)).$ (6)
@@ 4.3 Evaluation Process
4.4 Baselines
4.5 Implementation Settings
Table 1 compares four different evaluation approaches for computing BLEU-4 scores. The first approach uploads original article files to the GPT-4o playground without additional processing. The second approach calls the GPT-4o application programming interface (API) and integrates the user’s framework to process the files before uploading. The third approach applies the same logic but uses GPT-4o-mini, and the fourth approach also calls the GPT-4o-mini API with the user’s framework. This layout clarifies how different usage modes may affect the final BLEU-4 outcome. Tables 6, 7, 8, and 9 display the results from both using the GPT API and directly utilizing the GPT playground to analyze the article. In the GPT playground, the full episode script is uploaded, and the prompt ”Please evaluate episode $n$” is used to generate baseline results. When calling the GPT API, we upload the full episode and processed files, employ the OpenAI API combined with our framework to answer user queries, focusing primarily on evaluating the coherence of each episode. As noted earlier, the evaluation process has been optimized, resulting in a 90% improvement in analysis accuracy.
Table 1: Comparisons of four experimental settings for BLEU-4, contrasting direct file uploads with API-based approaches under both GPT-4o and GPT-4o-mini.
Evaluation Method BLEU-4
Table 2: Performance Comparison of LLMs with/without SCORE Framework
Model Consistency Coherence Item Status Complex QA
GPT-4 83.21 84.32 0 82.34
SCORE 85.61 ($\uparrow$2.4) 86.9 ($\uparrow$2.58) 98 ($\uparrow$98) 89.45 ($\uparrow$7.11)
GPT-4o 86.78 82.21 0 76.32
SCORE 88.68 ($\uparrow$1.9) 89.91 ($\uparrow$7.7) 96 ($\uparrow$96) 88.75 ($\uparrow$12.43)
Claude3 84.6 80.9 0 69.45
SCORE 87.2 ($\uparrow$2.6) 85.7 ($\uparrow$4.8) 93.1 ($\uparrow$93.1) 83.9 ($\uparrow$14.45)
Gemini-Pro 82.2 83.4 0 78.40
SCORE 85.2 ($\uparrow$3.2) 86.0 ($\uparrow$3.0) 95.0 ($\uparrow$95.0) 85.5 ($\uparrow$7.10)
Qwen-14B 79.3 76.2 0 25.15
SCORE 84.5 ($\uparrow$5.2) 79.1 ($\uparrow$2.9) 86.3 ($\uparrow$86.3) 61.2 ($\uparrow$36.05)
GPT-4o-mini 77.5 75.8 0 23.40
SCORE 81.9 ($\uparrow$4.4) 76.6 ($\uparrow$0.8) 79.8 ($\uparrow$79.8) 60.1 ($\uparrow$36.7)
Llama-13B 71.3 69.8 0 18.72
SCORE 79.1 ($\uparrow$7.8) 73.4 ($\uparrow$3.6) 76.2 ($\uparrow$76.2) 57.3 ($\uparrow$38.58)
4.6 Main Results
@@
As shown in Table 2, our experiments demonstrated that the proposed framework significantly improved the detection of narrative inconsistencies. Evaluations using the framework were able to more accurately detect inconsistencies in character actions or plot progression. The retrieval-augmented generation process helped GPT better filter irrelevant information, understand the current story context, and improved its ability to detect narrative continuity across multiple episodes. When quantitatively compared to baseline methods, such as using GPT model alone, the proposed framework showed substantial improvements in evaluation accuracy.
4.7 Case study
In this study, we generated several dozen stories using our framework and then performed evaluations on these outputs. We applied key word and summary extraction to each episode, with an average runtime of 13.233 seconds per episode (sampled from 10 episodes), and observed that each episode contained an average of 555 tokens. The evaluation was conducted using four metrics: Character Consistency, Plot Progression, Emotional and Psychological Realism, and Continuity and Consistency in Story Elements. The results indicate that, across both individual and combined evaluation metrics, most scores consistently fell in the range of three to four out of five. This finding suggests that, the model generally produced cohesive narratives. The results are shown in Figure 3.
Figure 3: Scores for our generated story on Character Consistency, Plot Progression, Emotional and Psychological Realism, and Continuity, and the overall score.
4.8 Ablation Studies
Ablation of LLM type. The comprehensive evaluation in Table 2 demonstrates SCORE’s consistent effectiveness across diverse model architectures, with particularly notable improvements for open-source models. The framework elevates baseline performance across all key metrics: consistency improvements range from 1.9% (GPT-4o) to 7.8% (Llama-13B), while coherence gains vary between 0.8% (GPT-4o-mini) and 7.7% (GPT-4o). Most dramatically, SCORE enables all models to achieve near-perfect item status recognition (79.8%-98%), where baseline performance was previously nonexistent. The results reveal an inverse relationship between model size and improvement magnitude - while commercial models like GPT-4 and Claude 3 show respectable gains of 2.4-4.8% in core metrics, smaller open-source models demonstrate more substantial improvements. Qwen-14B achieves 5.2% higher consistency and 2.9% better coherence, while Llama-13B shows the most dramatic consistency improvement at 7.8%. SCORE’s most transformative impact appears in complex QA tasks, where open-source models realize particularly significant gains: Qwen-14B improves by 36.05 percentage points (from 25.15% to 61.2%), GPT-4o-mini by 36.7 points (23.4% to 60.1%), and Llama-13B by 38.58 points (18.72% to 57.3%). These improvements substantially narrow the performance gap with commercial models, with Claude 3’s complex QA performance improving by 14.45 points (69.45% to 83.9%) and GPT-4’s by 7.11 points (82.34% to 89.45%). Notably, SCORE enhances GPT-4o’s performance beyond its successor GPT-4 in key metrics, improving coherence by 7.7 percentage points (from 82.21% to 89.91%) while GPT-4’s coherence improves by just 2.58 points. This demonstrates SCORE’s ability to extract additional capability even from state-of-the-art commercial models.
Table 3: Ablation results for our complete configuration (Full SCORE) versus removing each module.
Configuration Consistency Coherence Item Acc. Complex QA
Full SCORE 89.7 91.2 98.3 88.5
w/o Dynamic Tracking 72.1 ($\downarrow$17.6) 85.4 ($\downarrow$5.8) 61.2 ($\downarrow$37.1) 67.3 ($\downarrow$21.2)
w/o Context Summary 83.2 ($\downarrow$6.5) 68.7 ($\downarrow$22.5) 89.1 ($\downarrow$9.2) 71.4 ($\downarrow$17.1)
w/o Hybrid Retrieval 86.4 ($\downarrow$3.3) 88.9 ($\downarrow$2.3) 94.2 ($\downarrow$4.1) 77.6 ($\downarrow$10.9)
w/o Sentiment 87.1 ($\downarrow$2.6) 89.5 ($\downarrow$1.7) 96.8 ($\downarrow$1.5) 82.3 ($\downarrow$6.2)
Configurations analysis. Table 3 presents the performance of the full SCORE configuration alongside configurations with individual modules removed. As observed, removing the Dynamic Tracking causes a significant drop in both Consistency (from 89.7% to 72.1%, a decrease of 17.6 percentage points) and Item Accuracy (from 98.3% to 61.2%, a decrease of 37.1 percentage points), highlighting its critical role in maintaining narrative continuity and ensuring key item tracking. Similarly, excluding the Context Summary leads to a 22.5% drop in Coherence, indicating that summarizing context is vital for preserving the logical flow of the narrative. In contrast, the removal of the Hybrid Retrieval and Sentiment results in smaller performance declines, suggesting that while these components contribute positively, their impact is less pronounced compared to Dynamic Tracking and Context Summary.
Table 4: Multi-metric evaluation of SCORE’s robustness on GPT-4 using GPT-Story Dataset under various noise conditions
Condition Consistency Coherence Complex QA
Clean (noise-free) 92.1 89.7 88.9
30% Segment replacement 87.6 ($\downarrow$4.5) 85.2 ($\downarrow$4.5) 83.1 ($\downarrow$5.8)
$\sigma=0.3$ Similarity perturbation 85.9 ($\downarrow$6.2) 83.6 ($\downarrow$6.1) 80.7 ($\downarrow$8.2)
25% Sentiment reversal 83.3 ($\downarrow$8.8) 81.4 ($\downarrow$8.3) 78.2 ($\downarrow$10.7)
Combined noise 81.0 ($\downarrow$11.1) 79.1 ($\downarrow$10.6) 75.6 ($\downarrow$13.3)
Robustness Under Noisy Retrieval. Table 4 demonstrates SCORE’s remarkable noise resilience in the GPT-Story benchmark, showing controlled performance degradation across all test conditions. Under isolated noise scenarios (30% content replacement, $\sigma=0.3$ similarity distortion, or 25% sentiment inversion), key metrics exhibit limited degradation of just 4.5-10.7%. The framework shows particular sensitivity to emotional perturbations, with Complex QA performance dropping 10.7% under sentiment reversal - a finding that underscores the critical importance of affective continuity in multi-hop reasoning tasks. Most impressively, when subjected to combined noise interference (simultaneous application of all three noise types), SCORE maintains strong operational integrity with all critical metrics staying above 75%, despite a 10.6-13.3% overall performance reduction. These results provide conclusive evidence for the architectural robustness of SCORE’s hybrid design.

### v2

@@ Abstract
Large Language Models (LLMs) can generate creative and engaging narratives from user-specified input, but maintaining coherence and emotional depth throughout these AI-generated stories remains a challenge. In this work, we propose SCORE, a framework for Story Coherence and Retrieval Enhancement, designed to detect and resolve narrative inconsistencies. By tracking key item statuses and generating episode summaries, SCORE uses a Retrieval-Augmented Generation (RAG) approach, incorporating TF-IDF Ramos et al. (2003) and cosine similarity Rahutomo et al. (2012) to identify related episodes and enhance the overall story structure. Results from testing multiple LLM-generated stories demonstrate that SCORE significantly improves the consistency and stability of narrative coherence compared to baseline GPT models, providing a more robust method for evaluating and refining AI-generated narratives.
@@ 1 Introduction
Deep learning has transformed multiple domains including NLP, time series analysis and computer vision Qiu et al. (2025a, b, 2024); Li et al. (2024b); Li and Deng (2023); Li et al. (2024a). Large Language Models (LLMs) have demonstrated significant capabilities in generating long-form narratives, such as serialized stories or novels, by leveraging large-scale architectures and vast amounts of training data  Tao et al. (2024). However, maintaining narrative consistency over extended texts, especially in terms of character development and emotional coherence, remains a major challenge McAdams (2006). For instance, Khatun and Brown (2024) pointed out that achieving thematic consistency and managing dynamic plot states is crucial for maintaining the logical flow of a story. In practice, LLMs often struggle with inconsistencies when characters or key plot items reappear without proper explanation, disrupting the overall narrative structure.
Similarly, Liu et al. (2024) highlight the difficulties in managing multimodal elements within long-form narratives, noting that inconsistencies in character behavior or emotional tone can negatively impact reader engagement. These challenges indicate a need for more structured approaches in narrative generation that can better manage character arcs, plot developments, and emotional progression throughout the story.
In addition, recent works have highlighted the importance of memory mechanisms in LLM-based agents. Zhang et al. (2024) conducted a comprehensive survey on these mechanisms, identifying effective memory designs that help mitigate inconsistencies in narrative development—a challenge common to both interactive agents and narrative generation tasks. Additionally, Park et al. (2023) introduced generative agents that simulate human-like behavior using memory modules. These agents track the state of a wide array of interactable objects in a sandbox environment, ensuring consistent reasoning and enabling the smooth functioning of a simulated society. These researches inspired the design of our new framework.
In this work, we build upon recent advancements in Retrieval-Augmented Generation (RAG) Lewis et al. (2020), which dynamically incorporates relevant context to enhance narrative coherence. Expanding on these developments, we propose SCORE, a framework designed to evaluate three critical aspects of long-form narrative generation: character consistency, emotional coherence, and logical tracking of key plot elements. Our key contributions are:
@@
  • We outperform baseline GPT model Radford et al. (2019) in detecting continuity errors and maintaining overall narrative coherence.
@@ 2 Method
Our proposed method, SCORE, consists of three main components: (1) an LLM-based evaluation framework to assess the coherence of key story elements, (2) automatic generation of episode summaries to track plot development, and (3) a retrieval-augmented generation (RAG) approach that integrates the first two components, enabling enhanced user interaction and ensuring narrative consistency.
@@
Figure 1: The SCORE framework for improving AI-generated story coherence. (a) Extracts key item statuses per episode. (b) Conducts detailed analysis and summaries of each episode. (c) Uses RAG to answer user queries and resolve narrative inconsistencies.
2.1 Continuity Analysis and Key Item Status Correction
@@
By extracting key parts of GPT-4’s analysis, we identify instances where an item reappears in later episodes after being marked as lost or destroyed, without narrative explanation. Let $S_{i}(t)$ represent the state of item $i$ at time $t$, where $S_{i}(t)\in\{\text{active},\text{lost},\text{destroyed}\}$. If item $i$ reappears at time $t_{k}$ with $S_{i}(t_{k})=\text{active}$ after being previously marked as $S_{i}(t_{k-1})\in\{\text{lost},\text{destroyed}\}$, we flag this as a continuity error. To maintain consistency, the state remains $S_{i}(t_{k-1})$, avoiding an incorrect update. This approach systematically corrects discrepancies in item states, ensuring that narrative continuity is preserved by preventing erroneous state transitions.
2.2 Key Item Interaction Analysis
@@
For each episode, we conduct a thorough evaluation by summarizing key plot points, character actions, and tracking interactions with important items. Let $A_{c}(t)$ represent the actions of character $c$ at time $t$, and let $I_{i}(t)$ denote interactions with key item $i$. The model generates summaries that encapsulate essential elements, including $A_{c}(t)$ (character actions), relationships, and emotional changes across the episode. It then analyzes the specific interactions $I_{i}(t)$ between characters and key items, documenting these for further analysis. This step aggregates relevant narrative information—combining episode summaries, key item interactions $I_{i}(t)$, and character actions $A_{c}(t)$—to facilitate more precise future retrieval. The approach simplifies subsequent analysis of plot and item continuity, reducing redundancy and improving efficiency.
2.3 Similarity-Based Episode Evaluation and Sentiment Analysis
@@
We integrate similarity-based retrieval and sentiment analysis to improve episode evaluation and answer complex queries. It begins by loading summaries, full episode content, and key item states from structured JSON files. The content is segmented into smaller chunks using a text segmenter and embedded into a vector space model using FAISS Douze et al. (2024) and OpenAI embeddings. This vector space enables efficient retrieval of similar episodes for user queries or specific episode analysis.
@@
For evaluation, the system retrieves relevant past episodes by calculating cosine similarity scores between the current episode or query and all other episodes in the vector space. Let $S(e_{c},e_{p})$ represent the similarity score between the current episode $e_{c}$ and a past episode $e_{p}$. The top $N$ episodes with the highest $S(e_{c},e_{p})$ scores are retrieved for further analysis, providing a relevant summary of episodes for evaluation or answering questions.
@@ 3.1 Dataset Preparation
We collected a diverse set of stories generated by various GPT models, covering a wide range of genres such as science fiction, drama, and fantasy. Each story was divided into multiple episodes, with each episode containing the raw outputs generated by the language models, such as dialogues between characters, descriptions of events, and narrative elements. This variety in genres and structures allowed us to thoroughly evaluate our framework’s performance across different storytelling formats.
3.2 Baselines
Table 1: Performance comparison of models with and without SCORE
Model Consistency Coherence Item Status Complex Question
GPT-4o-mini 78.2 76.7 0 24.56
+ SCORE 82.6 (↑4.4) 77.5 (↑0.8) 80.5 (↑80.5) 63.0 (↑38.44)
GPT-4o 86.78 82.21 0 76.32
+ SCORE 88.68 (↑1.9) 89.91 (↑7.7) 96 (↑96) 88.75 (↑12)
GPT-4 83.21 84.32 0 82.34
+ SCORE 85.61 (↑2.4) 86.9 (↑2.58) 98 (↑98) 89.45 (↑7.11)
@@ 3.3 Evaluation Process
3.4 Metrics
3.5 Results
@@
Our experiments demonstrated that the proposed framework significantly improved the detection of narrative inconsistencies. Evaluations using the framework were able to more accurately detect inconsistencies in character actions or plot progression. The retrieval-augmented generation process helped GPT better filter irrelevant information, understand the current story context, and improved its ability to detect narrative continuity across multiple episodes. When quantitatively compared to baseline methods, such as using GPT model alone, the proposed framework showed substantial improvements in evaluation accuracy.
4 Limitations
While our SCORE framework offers significant improvements, it also presents several limitations:
  • Dependence on Retrieval Accuracy: The effectiveness of SCORE relies on accurately retrieving relevant episodes. Errors in similarity calculation or sentiment analysis may lead to the exclusion of important contextual information, affecting the overall evaluation quality.
  • Challenges in Capturing Emotional Nuances: Sentiment analysis may not fully capture the complexity of emotions, potentially missing subtle shifts or deeper emotional layers within the narrative.
  • Resource Intensive: Despite optimizations, SCORE still requires considerable computational resources for similarity calculations and sentiment analysis, which can be limiting for large datasets.
  • Generalizability: SCORE has been primarily tested on GPT-4-generated stories, and its performance on narratives from other models or with different structures remains an open question.